The Southern Corporation manufactures a single product and has the following cost structure: Variable costs per unit: Production $ 35 Selling and administrative $ 14 Fixed costs per year: Production $ 175,560 Selling and administrative $ 140,450 Last year, 7,980 units were produced and 7,680 units were sold. There was no beginning inventory. The carrying value on the balance sheet of the ending inventory of finished goods under variable costing would be:

Multiple Choice

  • $6,600 less than under absorption costing.

  • $7,680 less than under absorption costing.

  • the same as absorption costing.

  • $7,680 greater than under absorption costing.
Answer #1

Fixed manufacturing overhead per unit = 17,560/7,980

= 22 per unit

Fixed manufacturing overhead deferred in ending inventory

= 22 * (7,980-7,680) = 6,600

Option A is the answer

$6,600 less than under absorption costing